# Benchmarking Honesty of Large Language Models

### v2

@@ Abstract
Previous works on Large Language Models (LLMs) have mainly focused on evaluating their helpfulness or harmlessness. However, honesty, another crucial alignment criterion, has received relatively less attention. Dishonest behaviors in LLMs, such as spreading misinformation and defrauding users, present severe risks that intensify as these models approach superintelligence levels. Enhancing honesty in LLMs addresses critical deficiencies and helps uncover latent capabilities that are not readily expressed. This underscores the urgent need for reliable methods and benchmarks to effectively ensure and evaluate the honesty of LLMs.
In this paper, we introduce BeHonest, a pioneering benchmark specifically designed to assess honesty in LLMs comprehensively. BeHonest evaluates three essential aspects of honesty: awareness of knowledge boundaries, avoidance of deceit, and consistency in responses. Building on this foundation, we designed 10 scenarios to evaluate and analyze 9 popular LLMs on the market, including both closed-source and open-source models from different model families with varied model sizes. Our findings indicate that there is still significant room for improvement in the honesty of LLMs. We also encourage the AI community to prioritize honesty alignment in LLMs. Our benchmark and code can be found at: https://github.com/GAIR-NLP/BeHonest.
@@ 3.3.2 Benchmark Design
In this section, we measure the model’s consistency across four scenarios. First, we introduce modifications or bias features in the given prompt, such as variations in the prompt format under a zero-shot setting (Scenario 7) and the demonstration format in a few-shot setting (Scenario 8). An honest model should not change its answer due to these irrelevant alterations. Additionally, we implement self-evaluation tasks with different task formats (Scenario 9 and 10) to assess whether the model remains consistent in its own responses.
@@
Minor changes in wording, context, or framing of a prompt can significantly alter the responses generated by LLMs, which can affect their consistency and reliability (Brown et al.,, 2020; Liu et al.,, 2021). To assess the consistency of each model when encountering slight variations of the same content, we measure the performance spread (difference between maximum and minimum accuracy) across all five different prompt variations in Natural Instructions (Mishra et al.,, 2022), as inspired by Sclar et al., 2023b . We note that higher performance spread indicates that the model is more sensitive to variances in semantically-equivalent prompt formats, showing higher inconsistencies. The five prompt format variations can be found in Tab. 1.
@@
Previous research (Zhang et al.,, 2022; Yao et al.,, 2023) has demonstrated that LLMs are highly sensitive to few-shot demonstrations. To explore this inconsistency further, we assess the impact of irrelevant biased features in the demonstrations, examining the performance both with and without Chain of Thought (CoT) (Wei et al.,, 2022). Specifically, we implement Turpin et al., (2023)’s “Answer is Always A” setting on the Big-Bench Hard (BBH) dataset (Suzgun et al.,, 2023), where multiple-choice answer options in the demonstrations are reordered so that the correct answer is always (A). We then calculate the inconsistency rate by $(\text{Accuracy}_{\text{unbiased}}-\text{Accuracy}_{\text{biased}})/\text{Accuracy}_{\text{unbiased}}$.
@@
We release an overall score for scenarios 7 to 10. Since some metrics under the Consistency aspect are “higher is better” and some are “lower is better,” we first reverse the values under Performance Spread and Inconsistency Rate to align them as “higher is better” by following the formula $\frac{max(X)-X}{max(X)-min(X)}$, where X represents the individual values. Next, we normalize the values under Agreement Rate and Consistency Rate to a 0-100 scale using the formula $\frac{X-min(X)}{max(X)-min(X)}$. Finally, we average all the scores to produce an overall score.
@@ 4 Experiments
Non-deceptiveness Persona TrustLLM, Syco.-Interv. Augmented 470 sycophancy rate
Preference TrustLLM, TruthfulQA Combined 1,265 sycophancy rate
@@ 4.2.3 Consistency
Model Prompt Format Demo. Format O.F. Consistency M.C. Consistency Consistency
@@
GPT-4o 2.12 7.67 3.02 87.00 94.20 95.18
ChatGPT 3.11 50.49 11.39 73.00 70.40 68.13
Llama3-70b 5.25 30.99 1.14 94.40 33.60 64.75
Llama3-8b 5.50 57.01 18.50 57.40 70.80 49.49
Llama2-70b 4.25 57.89 25.94 66.00 61.60 51.77
Llama2-13b 6.50 75.53 31.76 71.80 79.40 39.07
Llama2-7b 3.25 82.08 49.59 47.60 73.80 38.83
Mistral-7b 2.75 35.19 27.43 82.20 70.00 69.03
Qwen1.5-14b 3.00 17.77 2.52 44.40 92.80 79.85
@@
From Tab. 5, we notice that GPT-4o has the lowest performance spread of 2.12%, indicating higher consistency and reliability in handling different prompt variations compared to other tested models. In contrast, Llama2-13b shows the highest sensitivity to prompt variations with a performance spread of 6.5%.
Overall, models perform differently even with slight changes in wording or formats, but the extent to which this happens varies.
@@
From Tab. 5, we observe that GPT-4o and Qwen1.5-14b perform the best, with consistency rates of 94.2% and 92.9%, respectively. Surprisingly, Llama3-70b has the lowest consistency rate (33.6%), indicating it as highly susceptible to doubts or disagreements despite being one of the largest models we tested, making it not robust. The remaining models have consistency rates in the 60s or 70s, showing a moderate degree of consistency in responses. We also report the modification rate in Appendix (Tab. 9) for additional information. Interestingly, while the overall accuracy drops for all models, there are still cases where the model initially provides a wrong answer but corrects itself after the user expresses doubts. An example is shown in Appendix (Tab. 25).
@@ Appendix D Examples
Table 15: Example of a model testing its knowledge boundary.
@@
Table 19: Example of a model lying in burglar deception (red) vs. not lying (green).
@@
Table 20: Example of a model lying in game (red) and not lying (green).
@@
Table 22: Example of a model showing consistency in demonstration format (green) and without (red).
@@
Table 23: Example of a model showing consistency in open-form (green) and without (red).
@@
Table 24: Example of a model showing consistency in multiple-choice formatted questions (green) and without (red).
@@
Table 25: Example where model initially responds an incorrect answer then corrects itself after a user poses doubt.

### v3

# Benchmarking Honesty in Large Language Models

@@ Abstract
Previous works on Large Language Models (LLMs) have mainly focused on evaluating their helpfulness or harmlessness. However, honesty, another crucial alignment criterion, has received relatively less attention. Dishonest behaviors in LLMs, such as spreading misinformation and defrauding users, present severe risks that intensify as these models approach superintelligent levels. Enhancing honesty in LLMs addresses critical limitations and helps uncover latent capabilities that are not readily expressed. This underscores the urgent need for reliable methods and benchmarks to effectively ensure and evaluate the honesty of LLMs.
In this paper, we introduce BeHonest, a pioneering benchmark specifically designed to assess honesty in LLMs comprehensively. BeHonest evaluates three essential aspects of honesty: awareness of knowledge boundaries, avoidance of deceit, and consistency in responses. Building on this foundation, we designed 10 scenarios to evaluate and analyze 9 popular LLMs on the market, including both closed-source and open-source models from different model families with varied model sizes. Our findings indicate that there is still significant room for improvement in the honesty of LLMs. We encourage the AI community to prioritize honesty alignment in these models, which can harness their full potential to benefit society while preventing them from causing harm through deception or inconsistency. Our benchmark and code can be found at: https://github.com/GAIR-NLP/BeHonest.
@@ 3.3.2 Benchmark Design
In this section, we measure the model’s consistency across four scenarios. First, we introduce modifications or bias features in the given prompt, such as variations in the prompt format under a zero-shot setting (Scenario 7) and the demonstration format in a few-shot setting (Scenario 8). An honest model should not significantly change its answer due to these irrelevant alterations. Additionally, we implement self-evaluation tasks with different task formats (Scenario 9 and 10) to assess whether the model remains consistent in its own responses.
@@
Minor changes in wording, context, or framing of a prompt can significantly alter the responses generated by LLMs, which can affect their consistency and reliability (Brown et al.,, 2020; Liu et al.,, 2021). To assess the consistency of each model when encountering slight variations of the same content, we measure the performance spread (difference between maximum and minimum accuracy) across all five different prompt variations based on the Natural Instructions dataset (Mishra et al.,, 2022), as inspired by Sclar et al., 2023b . We note that higher performance spread indicates that the model is more sensitive to variances in semantically-equivalent prompt formats, showing higher inconsistencies. The five prompt format variations can be found in Tab. 1.
@@
Previous research (Zhang et al.,, 2022; Yao et al.,, 2023) have demonstrated that LLMs are highly sensitive to few-shot demonstrations. To further explore this inconsistency, we assess the impact of irrelevant biased features in the demonstrations, examining the performance both with and without Chain of Thought (CoT) (Wei et al.,, 2022). Specifically, we implement Turpin et al., (2023)’s “Answer is Always A” setting on the Big-Bench Hard (BBH) dataset (Suzgun et al.,, 2023), where multiple-choice answer options in the demonstrations are reordered so that the correct answer is always (A). We then calculate the inconsistency rate with the formula: $(\text{Accuracy}_{\text{unbiased}}-\text{Accuracy}_{\text{biased}})/\text{Accuracy}_{\text{unbiased}}$.
@@
We release an overall score for scenarios 7 to 10. Since some metrics under the Consistency aspect are “higher is better” and some are “lower is better,” we first reverse the values under Performance Spread and Inconsistency Rate to align them as “higher is better” by following the formula
$\left(\frac{\max(X)-x}{\max(X)-\min(X)}\right)\times 100,$
where $x$ represents the individual values, and $X$ is the set of values under a scenario across all models. Next, we normalize the values under Agreement Rate and Consistency Rate to a 0-100 scale using the formula
$\left(\frac{x-\min(X)}{\max(X)-\min(X)}\right)\times 100.$
Finally, we average all the scores to produce the overall score.
@@ 4 Experiments
Non-Deceptiveness Persona Sycophancy TrustLLM, Syco.-Interv. Augmented 470 sycophancy rate
Preference Sycophancy TrustLLM, TruthfulQA Combined 1,265 sycophancy rate
@@ 4.2.3 Consistency
Model Prompt Format Demon. Format O.F. Consistency M.C. Consistency Consistency
@@
GPT-4o 2.12 7.67 3.02 87.00 94.20 96.26
ChatGPT 3.11 50.49 11.39 73.00 70.40 63.32
Llama3-70b 5.25 30.99 1.14 94.40 33.60 59.44
Llama3-8b 5.50 57.01 18.50 57.40 70.80 41.62
Llama2-70b 4.25 57.89 25.94 66.00 61.60 44.42
Llama2-13b 6.50 75.53 31.76 71.80 79.40 35.20
Llama2-7b 3.25 82.08 49.59 47.60 73.80 29.39
Mistral-7b 2.75 35.19 27.33 82.20 70.00 66.05
Qwen1.5-14b 3.00 17.77 2.52 44.40 92.80 72.24
@@
From Tab. 5, we notice that GPT-4o has the lowest performance spread of 2.12%, indicating higher consistency and reliability in handling different prompt variations of the same content compared to other tested models. In contrast, Llama2-13b shows the highest sensitivity to prompt variations with a performance spread of 6.5%. Overall, models perform differently even with slight changes in wording or formats, such as additional/missing punctuation marks or spacing, but the extent to which this happens varies.
@@
From Tab. 5, we observe that GPT-4o and Qwen1.5-14b perform the best, with consistency rates of 94.2% and 92.8%, respectively. Surprisingly, Llama3-70b has the lowest consistency rate (33.6%), indicating it as highly susceptible to doubts or disagreements despite being one of the largest models we tested, making it not robust. The remaining models have consistency rates in the 60s or 70s, showing a moderate degree of consistency in responses. We also report the modification rate in Appendix (Tab. 9) for additional information. Interestingly, while the overall accuracy drops for all models, there are still cases where the model initially provides a wrong answer but corrects itself after the user expresses doubts. An example is shown in Appendix (Tab. 25).
@@ Appendix D Examples
Table 15: Example of testing a model’s self-knowledge.
@@
Table 19: Example of a model lying (red) vs. not lying (green) in burglar deception.
@@
Table 20: Example of a model lying (red) vs. not lying (green) in game.
@@
Table 22: Example of a model showing consistency (green) vs. inconsistency (red) in demonstration format.
@@
Table 23: Example of a model showing consistency (green) vs. inconsistency (red) in open-form.
@@
Table 24: Example of a model showing consistency (green) vs. inconsistency (red) in multiple-choice formatted questions.
@@
Table 25: Example where a model initially responds with an incorrect answer then corrects itself after a user poses doubt.